# Parameter-Efficient Fine-Tuning of Large Language Models for Unit Test Generation: An Empirical Study

### v1

@@ IV-B Fine-tuning methods selection
In line with other empirical studies on PEFT [8, 31], we select the popularly used PEFT methods: LoRA [15], (IA)3 [16], and prompt tuning [17]. They are readily available in popular libraries and exemplify most of the primary approaches within PEFT.
(a) Diagram of LoRA (based on [15]).
(b) Diagram of (IA)3 (based on [16]).
(c) Diagram of prompt tuning (based on [17]).
@@
  • LoRA: A reparametrization-based method that updates a weight matrix by training a new matrix that is decomposed into a product of two matrices of lower rank: $\delta W=W_{A}\cdot W_{B}$, where $W_{A}\in\mathbb{R}^{\text{in}\times r}$ and $W_{B}\in\mathbb{R}^{r\times\text{out}}$ [15]. As demonstrated in Figure 1(a), the decomposed matrix is then added to the original frozen K and V projection matrices in the attention modules. All other modules are kept frozen.
@@
  • (IA)3: An additive method that focuses on fine-tuning specific intermediate activations within the model [16]. As illustrated in Figure 1(b), this is done by learning three new vectors: $l_{v},l_{k},l_{ff}$; rescaling key, value, and hidden feed-forward activations, respectively.
@@
  • Prompt tuning: A soft prompts method that fine-tunes a subset of the model’s input embeddings. Exemplified in Figure 1(c), the model input embeddings are prepended with a trainable tensor $P\in\mathbb{R}^{N\times D}$ (known as a “soft prompt”) [17]. Each element of $P$ is also referred to as a “virtual token.” During training, these “virtual tokens” are updated while all the rest of the model parameters are kept frozen.
IV-C Models
@@
For our experiments, we select a diverse set of open-source decoder-only model families trained on code to evaluate the performance and efficiency of PEFT methods. The chosen models span a range of sizes and architectures and are trained on diverse datasets and training objectives. This offers extensive insights into the interplay between various model capacities and PEFT methods. Further, all models are ensured to support the Java programming language to facilitate a fair comparison of our benchmarking datasets described in Section IV-D.
@@
  • Salesforce CodeGen Models: To democratize program synthesis, Salesforce AI Research developed CodeGen [32]. Pioneering in multi-turn program synthesis, CodeGen presents a variety of sizes, ranging from 350 million to 16 billion parameters. These models have undergone training across a diverse array of both programming and natural languages. The models were trained on The pile [33], followed by a code subset of Google’s BigQuery datasets Cloud [34]. An updated version CodeGen2 [35] is also available in sizes 1 billion to 16 billion, trained on The Stack dataset [36] using both causal language modeling (CLM) and file-level span corruption. We select the smallest model of the first generation CodeGen family (CodeGen-350M-multi), along with all model sizes of the second generation (1 billion, 3.7 billion, 7 billion, and 16 billion parameters).
@@
  • Meta LLaMA Models: Code Llama [37], derived from Llama 2 [38], is a specialized model further fine-tuned on the code-specific parts of its pre-training data. With enhanced coding capabilities, it can generate code, interpret natural language prompts, assist in code completion, and aid debugging. It comes in three variants. One is optimized for instructions. Another for Python. We select the base variant CodeLlama-7B, as it is designed for general code synthesis and understanding and supports a wide range of popular languages, including Java.
@@
  • BigCode StarCoder Models: This model family includes the 15.5 billion parameters StarCoderBase model [39], along with an updated model generation named StarCoder2 [40], available in 3 billion, 7 billion, and 15 billion parameters. StarCoderBase was trained on 80+ programming languages from The Stack dataset [36]. The updated StarCoder2 is trained on 619 programming languages and other data sources, such as GitHub pull requests, Kaggle notebooks, and code docs. We select all model generations and sizes in this model family.
IV-D Datasets
IV-D1 Methods2Test
@@
To ensure our experiments’ validity and practical applicability, we rely on Methods2Test [1]. It is the largest publicly available corpus of Java unit test cases and corresponding focal methods. This dataset contains 780,000 test cases sourced from 91,000 open-source repositories on GitHub. These real-world test cases represent the complexity and variability inherent in software development, making them an ideal choice for evaluating automated test generation methods. By leveraging Methods2Test, we can assess the performance of the various PEFT methods and models in generating high-quality unit tests that align with real-world coding practices and challenges.
@@
One of the advantages promised by various PEFT methods is the reduction in the amount of data needed for effective training. Thus, we opt to downsize the Methods2Test dataset to enhance efficiency without compromising the representativeness of the data. Specifically, we adopt a strategy wherein only one focal function is selected from each class in the dataset. By doing so, we maintain the essence and diversity of the original dataset while significantly reducing its size. Further, the Methods2Test dataset is distributed in deconstructed components. Hence, we assemble it according to the structure seen in LABEL:lst:data-structure and repackage it as Methods2Testsmall [^3] Available at https://huggingface.co/datasets/andstor/methods2test_small. The resulting dataset contains 7,44K training, 953 validation, and 1.02K testing tests.
@@
LISTING I: Dataset components: focal method (fm), focal class name (fc), constructor signatures (c), method signatures, fields (f), test case (t), test class name (tc). Data is assembeld according to fm+fc+c+m+f+t+tc. The code is split after the test case signature. Everything above the red line is the source input for unit test generation. The code below the red line is the target output of the unit test generation.
IV-D2 HumanEval-X
HumanEval is a benchmarking dataset for code generation, first introduced by Chen et al. [41]. It consists of 164 hand-written programming problems written in Python. For each programming problem, a function signature, docstring, body, and several unit tests are provided. Zheng et al. [42] extended the dataset to multiple languages, named HumanEval-X. We use the Java subset of the HumanEval-X dataset to evaluate potential adverse fine-tuning effects.
IV-E Training
As the training dataset Methods2Testsmall only includes method signatures for non-focal methods, the unit test generation task can be considered a sequence-to-sequence (seq2seq) problem. For training the decoder-only models described in Section IV-C, we, therefore, model the seq2seq task as an autoregressive problem. This is done by only calculating loss based on the target code, i.e., the unit test code. We minimize the following autoregressive cross-entropy loss function:
@@ IV-E Training
TABLE I: Model-agnostic hyperparameters for training.
@@
We use the Transformers [43] and PEFT libraries [44] from Hugging Face to conduct all training. We use NVIDIA A100 GPUs and leverage DeepSpeed ZeRO-Offload [45] where necessary. We use the hyperparameters shown in Table I for training. Like other studies, e.g., [8], we select hyperparameter values that align with those specified in the original papers of the respective PEFT techniques. Weyssow et al. [8] fine-tuned their models for up to five epochs. Our results show that the losses of most of our fine-tuned models are stable after three epochs [^4] Available at https://huggingface.co/andstor/peft-unit-test-generation-experiments. Regarding the unspecified hyperparameters, we leave them to library defaults. Following defaults of the PEFT library, rank decomposition for LoRA is applied to the attention layers. The same attention layers and the default feed-forward modules (see Table II) are targeted for (IA)3. The selected models have context windows ranging from 2048 to 16,384. To make the experiments align with each other, we limit the number of tokens to the smallest window of 2048.
TABLE II: Model-specific hyperparameters for training.
@@
Targeted attention modules LoRA, (IA)3 codegen-350M-multi qkv_proj
@@
Targeted feedforward modules (IA)3 codegen-350M-multi fc_out
@@
IV-F Generation
For generating, we provide the source part of the data (see LABEL:lst:data-structure) as input to the models, while the target part (original code) is used as the ground truth. The generated unit test code is then compared to the ground truth to calculate their similarities. We implement a stopping strategy based on matching braces to generate well-formed functions. For incomplete unit tests, we follow the protocol from [2] where we iteratively delete lines (from bottom to top), add one to two curly brackets, and check if the syntax is valid using an ANTLR v4-based [46] Java parser [47]. We only evaluate samples that are well-formed.
We also apply DeepSpeed-Inference [48] to facilitate model parallel (MP) to fit the larger models that would otherwise not fit in GPU memory. Table III shows the hyperparameter details used for generation.
TABLE III: Hyperparameters for generation.
@@ IV-F Generation
IV-G Metrics
IV-G1 Syntactic Correctness
@@
Following Yuan et al. [3], we measure the syntactic correctness of the generated unit tests, ensuring that they conform to the grammar of the Java programming language. After the fixing described in Section IV-F, we employ the same parser as described in Section IV-F. This involves parsing the generated tests and verifying that they are free of syntax errors. We then report the percentage of unit tests with valid syntax for each model.
IV-G2 Similarity
@@
BLEU score [49] is a very common metric used to evaluate natural language machine translation. It measures the similarity of the machine-translated text to the reference translations. However, Ren et al. [50] recognized that it neglects important syntactic and semantic features of programming languages. To remedy this, they developed CodeBLEU [50]. Based on n-gram matching used in BLEU, they further augment the scoring by leveraging abstract syntax trees (AST) and code semantics via data flow. To evaluate the primary performance of the various PEFT methods, we use CodeBLEU to measure the similarity between the generated unit test code and the ground truth.
IV-G3 Coverage
We also execute the generated tests for the HumanEval-Xjava, and calculate code coverage. Specifically, we calculate the pass@1 metric [41] and instruction and branch coverage using the Java Code Coverage Library JaCoCo [51]. Instruction coverage measures the percentage of executed instructions, while branch coverage evaluates the percentage of executed control flow branches in the program.
V Experimental Results

### v2

@@ 4.2 Fine-tuning methods selection
In line with other empirical studies on PEFT [weyssow2024exploring, Pu2023EmpiricalAO], we select the most popularly used PEFT methods: LoRA [Yu2023LowRankAO], (IA)3 [Liu2022FewShotPF], and prompt tuning [Lester2021ThePO]. They are readily available in popular libraries and exemplify most of the primary approaches within PEFT.
(a) Diagram of LoRA (based on [Yu2023LowRankAO]).
(b) Diagram of (IA)3 (based on [Liu2022FewShotPF]).
(c) Diagram of prompt tuning (based on [Lester2021ThePO]).
@@
  • LoRA: A reparametrization-based method that updates a weight matrix by training a new matrix that is decomposed into a product of two matrices of lower rank: $\delta W=W_{A}\cdot W_{B}$, where $W_{A}\in\mathbb{R}^{\text{in}\times r}$ and $W_{B}\in\mathbb{R}^{r\times\text{out}}$ [Yu2023LowRankAO]. As demonstrated in Figure 1(a), the decomposed matrix is then added to the original frozen K and V projection matrices in the attention modules. All other modules are kept frozen.
@@
  • (IA)3: An additive method that focuses on fine-tuning specific intermediate activations within the model [Liu2022FewShotPF]. As illustrated in Figure 1(b), this is done by learning three new vectors: $l_{v},l_{k},l_{ff}$; rescaling key, value, and hidden feed-forward activations, respectively.
@@
  • Prompt tuning: A soft prompts (additive) method that fine-tunes a subset of the model’s input embeddings. Exemplified in Figure 1(c), the model input embeddings are prepended with a trainable tensor $P\in\mathbb{R}^{N\times D}$ (known as a “soft prompt”) [Lester2021ThePO]. Each element of $P$ is also referred to as a “virtual token.” During training, these “virtual tokens” are updated while all the rest of the model parameters are kept frozen.
4.3 Models
@@
We select a diverse set of open-source autoregressive model families trained on code to evaluate the performance and cost-efficiency of PEFT methods. The chosen models span a range of sizes and architectures from the last four years, trained on diverse datasets and training objectives. This offers extensive insights into the interplay between various model capacities and PEFT methods. Further, all models are ensured to support the Java programming language to facilitate a fair comparison of our benchmarking datasets described in Section 4.4.
@@
  • Salesforce CodeGen Models: To democratize program synthesis, Salesforce AI Research developed CodeGen [nijkamp2022codegen]. Pioneering in multi-turn program synthesis, CodeGen presents a variety of sizes, ranging from 350 million to 16 billion parameters. These models have undergone training across a diverse array of both programming and natural languages. The models were trained on “The pile” [gao2020pile], followed by a code subset of Google’s BigQuery datasets [GoogleBigQuery]. An updated version CodeGen2 [nijkamp2023codegen2] is also available in sizes 1 billion to 16 billion, trained on “The Stack” dataset [Kocetkov2022TheStack] using both causal language modeling (CLM) and file-level span corruption. We select the smallest model of the first generation CodeGen family (CodeGen-350M-multi), along with all model sizes of the second generation (1 billion, 3.7 billion, 7 billion, and 16 billion parameters).
@@
  • Meta LLaMA Models: Code Llama [Rozire2023CodeLO], derived from Llama 2 [Touvron2023Llama2O], is a specialized model further fine-tuned on the code-specific parts of its pre-training data. With enhanced coding capabilities, it can generate code, interpret natural language prompts, assist in code completion, and aid debugging. It comes in three variants. One is optimized for instructions. Another for Python. We select the base variant CodeLlama-7B, as it is designed for general code synthesis and understanding and supports a wide range of popular languages, including Java.
@@
  • BigCode StarCoder Models: This model family includes the 15.5 billion parameters StarCoderBase model [li2023starcoder], along with an updated model generation named StarCoder2 [Lozhkov2024StarCoder2A], available in 3 billion, 7 billion, and 15 billion parameters. StarCoderBase was trained on 80+ programming languages from “The Stack” dataset [Kocetkov2022TheStack]. The updated StarCoder2 is trained on 619 programming languages and other data sources, such as GitHub pull requests, Kaggle notebooks, and code docs. We select all model generations and sizes in this model family.
  • Qwen 2.5 Coder Models: Qwen 2.5 Coder [hui2024qwen] is the latest code-specific series of Qwen LLMs. The series comes in 0.5, 1.5, 3, 7, 14, 32 billion parameters. The models are built upon the Qwen2.5 architecture and pre- trained on a vast corpus of over 5.5 trillion tokens. The dataset is sourced from a wide range of public code repositories, as well as large-scale web-crawled data containing code-related texts. We select the 3, 7 and 14 bllion parameter sizes of this model family.
Table 1 shows model statistics for the selected models derived from cfahlgren1/hub-stats hugging face dataset [^3] https://huggingface.co/datasets/cfahlgren1/hub-stats as of 13 November 2025. The chosen models span a representative set of model families released over the past four years, many of which remain highly popular.
Table 1: HuggingFace model download statistics from cfahlgren1/hub-stats dataset as of 13 November 2025.
Model Downloads in October 2025 Downloads all time Created at
Qwen/Qwen2.5-Coder-14B 4621 185216 2024-11-08
Qwen/Qwen2.5-Coder-3B 15469 213837 2024-11-08
Qwen/Qwen2.5-Coder-7B 22839 453481 2024-09-16
meta-llama/CodeLlama-7b-hf 2274 200519 2024-03-13
bigcode/starcoder2-7b 17013 695181 2024-02-20
bigcode/starcoder2-15b 5397 450456 2024-02-20
bigcode/starcoder2-3b 192177 6168133 2023-11-29
bigcode/starcoderbase 7185 210970 2023-05-03
Salesforce/codegen2-16B_P 13 37908 2023-04-26
Salesforce/codegen2-7B_P 90 28639 2023-04-26
Salesforce/codegen2-3_7B_P 42 39302 2023-04-25
Salesforce/codegen2-1B_P 226 77365 2023-04-25
Salesforce/codegen-350M-multi 4113 588888 2022-04-11
4.4 Datasets
4.4.1 Methods2Testsmall
@@
To ensure our experiments’ validity and practical applicability, we rely on Methods2Test by [tufano2021unit]. It is the largest publicly available corpus of Java unit test cases and corresponding focal methods. This dataset contains 780,000 test cases sourced from 91,000 open-source repositories on GitHub. These real-world test cases represent the complexity and variability inherent in software development, making them an ideal choice for evaluating automated test generation methods. By leveraging Methods2Test, we can assess the performance of the various PEFT methods and models in generating high-quality unit tests that align with real-world coding practices and challenges.
@@
One of the advantages promised by various PEFT methods is the reduction in the amount of data needed for effective training. Thus, we opt to downsize the Methods2Test dataset to enhance efficiency without compromising the representativeness of the data. Specifically, we adopt a strategy wherein only one focal function is selected at random from each class in the dataset. By doing so, we maintain the essence and diversity of the original dataset while significantly reduce its size. Further, the Methods2Test dataset is distributed in deconstructed components. Hence, we assemble it according to the structure seen in LABEL:lst:data-structure and package it as Methods2Testsmall. The resulting dataset contains 7440 training, 953 validation, and 1017 testing tests. We use this as the primary training dataset.
4.4.2 Methods2Testrunnable
As the Methods2Test dataset is made up of tests from the real world, executing the tests to measure their passing rates, test coverages, and mutation scores needs significant extra work. This includes retrieving the correct repository version, resolving dependencies, injecting coverage collection tools and mutation testing systems into the existing build system, and finally executing the code in a controlled and safe manner. We construct a runnable subset of Methods2Test by the following steps: 1) Golden commit: Methods2Test does not provide the repository commits at the time of scraping. Therefore, for each repository, we consider all commits prior to the last commit [^4] https://github.com/microsoft/methods2test/commit/309ee31cb3752cd8f3ad4b33cfd2b6e5b0ec5029 on GitHub that updated the Methods2Test dataset (2021-05-18T23:18:51.000Z UTC) in which the focal method and test class files exist as candidate commits. We brute force (begin at the most recent candidate commit and iteratively check older commits) compare all available dataset data against the repository until we find an exact match—a golden commit. 2) Build system: We target repositories using the Maven build system [apache-maven:3.8.1], as this is somewhat structured and not as flexible as Gradle. We rely on build file names to determine the build system. Of the 1017 tests in the test split of Methods2Testsmall, 624 of the repositories in Methods2Test are using Maven, 264 are using Gradle, and 129 are using others. 3) Runnable: Each repository in Methods2Test’s test split is then checked for buildability by running Maven and executing the test split of Methods2Testsmall. Due to a lot of old and missing dependencies, only 119 of the 624 Maven repositories are buildable. We collect all tests from Methods2Test with buildable repositories, resulting in 6150 tests. We then validate their runnability by executing every test, obtaining 3712 runnable tests (see section 4.8 for additional test execution details.) 4) Assertion heuristic: Finally, we filter out any tests that do not contain any of the substrings “assert”, “verify”, or “fail”. These substrings are common indicators of meaningful test logic in Java-based testing frameworks. Tests lacking these keywords are often empty, incomplete, disabled, or rely on custom test runners and frameworks that bypass standard test assertions. To avoid misleading results—such as treating a test that silently does nothing as correct—we conservatively discard such test cases. This results in a total of 3266 runnable tests. We package it as Methods2Testrunnable. We use this as the primary testing dataset. We package the commit, build system, and execution status as Methods2Testmeta.
@@ 4.4.2 Methods2Testrunnable
Listing 1: Dataset components: focal method (fm), focal class name (fc), constructor signatures (c), method signatures(m), fields (f), test case (t), test class name (tc). Data is assembeld according to fm+fc+c+m+f+t+tc. The code is split after the test case signature. Everything above the red line is the source input for unit test generation. The code below the red line is the target output of the unit test generation.
4.4.3 HumanEval-X
HumanEval is a benchmarking dataset for code generation, first introduced by [Chen2021EvaluatingLL]. It consists of 164 hand-written programming problems written in Python. For each programming problem, a function signature, docstring, body, and several unit tests are provided. [Zheng2023CodeGeeXAP] extended the dataset to multiple languages, named HumanEval-X. We use the Java subset of the HumanEval-X dataset to evaluate potential catastrophic forgetting effects of fine-tuning.
4.5 Training
As the training dataset Methods2Testsmall only includes method signatures for non-focal methods, the unit test generation task can be considered a sequence-to-sequence (seq2seq) problem. Thus, to preserve the code completion capabilities of the decoder-only models described in Section 4.3, we model the seq2seq task as an autoregressive problem. This is done by only calculating loss based on the target code, i.e., the unit test code. We minimize the following autoregressive cross-entropy loss function:
@@ 4.5 Training
Table 2: Model-agnostic hyperparameters for training.
@@
Table 3: Model-specific hyperparameters for training.
@@
Targeted attention modules LoRA, (IA)3 Salesforce/codegen-350M-multi qkv_proj
@@
Qwen/Qwen2.5-Coder-3B q_proj,v_proj
Qwen/Qwen2.5-Coder-7B q_proj,v_proj
Qwen/Qwen2.5-Coder-14B q_proj,v_proj
Targeted feedforward modules (IA)3 Salesforce/codegen-350M-multi fc_out
@@
Qwen/Qwen2.5-Coder-3B down_proj
Qwen/Qwen2.5-Coder-7B down_proj
Qwen/Qwen2.5-Coder-14B down_proj
We use the Transformers [wolf-etal-2020-transformers] and PEFT libraries [peft] from Hugging Face to conduct all training. We use NVIDIA A100 GPUs and leverage DeepSpeed ZeRO-Offload [Ren2021ZeROOffloadDB] where necessary. We use the hyperparameters shown in Table 2 for training. Like other studies, e.g. that of [weyssow2024exploring], we select hyperparameter values that align with those specified in the original papers of the respective PEFT techniques. [weyssow2024exploring] fine-tuned their models for up to five epochs. Our results show that the losses of most of our fine-tuned models are fairly stable after three epochs [^5] Available at https://huggingface.co/datasets/andstor/peft-unit-test-generation-experiments. Regarding the unspecified hyperparameters, we leave them to library defaults. Following the defaults of the PEFT library, rank decomposition for LoRA is applied to the attention layers. The same attention layers and the default feed-forward modules are targeted for (IA)3. An exhaustive set of selected hyperparameters is provided [^6] https://huggingface.co/datasets/andstor/peft-unit-test-generation-experiments#training-hyperparameters for reproducing our results. The selected models have context windows ranging from 2048 to 16,384. To make the experiments align with each other, we limit the number of tokens to the smallest window of 2048.
4.6 Generation
Table 4: Hyperparameters for generation.
@@ 4.6 Generation
For generating unit tests with the Methods2Testrunnable dataset, we provide the source part of the data (exemplified in LABEL:lst:data-structure) as input to the models to generate unit test cases. In order to facilitate reliable execution during later evaluation stages, the source also includes the test method function definition. For the evaluation of catastrophic forgetting, we use the default evaluation protocol for the HumanEval-X dataset. This includes generating the body of a focal method, given an function signature and docstring as the source input.
As we use a context window of 2048, we filter out any samples where the source + target overflows. We implement a stopping strategy based on matching braces to generate well-formed functions. That is, functions that have matching start and end braces. For incomplete unit tests, we follow the protocol from [siddiq2024using] where we iteratively delete lines (from bottom to top), add one to two curly brackets, and check if the syntax is valid using an ANTLR v4-based [antlr4] Java parser [java-antlr4]. We only evaluate samples that are well-formed. We use greedy decoding without any sampling or frequency penalty. Table 4 shows the hyperparameter details used for generation.
4.7 Metrics
4.7.1 Syntactic Correctness
@@
Following [yuan2023manual], we measure the syntactic correctness of the generated unit tests, ensuring that they conform to the grammar of the Java programming language. After the fixing described in Section 4.6, we employ the same parser as described in Section 4.6. This involves parsing the generated tests and verifying that they are free of syntax errors. We then report the percentage of unit tests with valid syntax for each model.
4.7.2 Similarity
@@
BLEU score by [Papineni2002BleuAM] is a very common metric used to evaluate natural language machine translation. It measures the similarity of the machine-translated text to the reference translations. However, [Ren2020CodeBLEUAM] recognized that the BLEU score neglects important syntactic and semantic features of programming languages. To remedy this, they developed CodeBLEU [Ren2020CodeBLEUAM]. Based on n-gram matching used in BLEU, they further augment the scoring by leveraging abstract syntax trees (AST) and code semantics via data flow. To evaluate the primary performance of the various PEFT methods, we use CodeBLEU to measure the similarity between the generated unit test code and the ground truth. In our experiments, we use the tokenizer_13a  [^7] https://github.com/mjpost/sacrebleu/blob/master/sacrebleu/tokenizers/tokenizer_13a.py from SacreBLEU by [bojar2018call], equivalent to the standard mteval-v13a tokenizer used by the Conference on Machine Translation (WMT).
4.7.3 Passing Rate
We use pass@1, a special case of the pass@k metric introduced by [Chen2021EvaluatingLL], to assess the correctness of generated outputs. In our setting, given a single focal method and one corresponding generated test case, pass@1 measures the probability that the test passes when executed. This reflects a realistic deployment scenario where there is only one chance to provide the user with a useful generated test case. Since our generation does not involve sampling, this also ensures reproducibility.
4.7.4 Coverage
To evaluate the effectiveness of the tests generated using Methods2Testrunnable, we calculate the coverage of the tests that pass. Specifically, we measure instruction coverage and branch coverage. These metrics provide a good indication of how thoroughly the test cases exercise the method under test.
4.7.5 Mutation Score
We evaluating the semantic strength of the generated tests by conducting mutation testing of the . Mutation testing evaluates the fault-detection capability of test cases by introducing small artificial defects, so-called mutants, into the program under test and checking whether the generated tests can detect them. A mutant is considered killed if the test fails when the mutation is present. We compute the mutation score as the number of mutants killed divided by the total number of mutants, following standard practice in software testing research [jia2011an].
4.8 Test Execution
To safely execute untrusted code, we utilize Docker containers that isolate the test environment. Java tests from our datasets are executed using the Apache Maven build system [apache-maven:3.8.1]. For the Methods2Testrunnable dataset, the Maven configuration is provided by cloning the test’s repository. For HumanEval-X, we use a hardcoded setup of Maven, and write the test and focal method to files. To measure code coverage, we employ the JaCoCo Java Code Coverage Library [jacoco:0.8.3]. For conducting mutation testing, we use the JVM mutation testing system Pitest [pitest:1.16.0]. Pitest inserts mutants into jvm by re-writing the class after it has loaded. We use the following default mutators: “Conditionals Boundary”, “Increments”, “Invert Negatives”, “Mat”, “Negate Conditionals”, “Void Method Calls”, “Empty returns”, “False Returns”, “True returns”, “Null returns”, and “Primitive return”.
Prior to execution, we replace the original implementation with the generated content. In the case of Methods2Testrunnable, this involves replacing the test method. For HumanEval-X, it involves replacing the focal method. We then instruct Maven to execute the single corresponding unit test method and rely on Maven to retrieve any dependencies and build required modules.
Test execution results are obtained from the XML output produced by the standard Maven Surefire Plugin [mavensurefireplugin]. A test run may conclude in several ways. It may “succeed” or “fail” based on assertion outcomes. It may also result in an error. Errors typically arise from unexpected runtime issues, such as unhandled exceptions during execution—an “interruption”. Beyond execution-specific issues, the environment can also lead to problems. These include “compilation errors”—often due to missing references, type mismatches, or unfulfilled dependencies. Although our datasets generally exclude unbuildable codebases (see section 4.4.2), compilation errors can still occur with generated content. Other environmental exceptions may arise during container setup, file I/O, or subprocess invocation, causing some tests to fail to run.
5 Experimental Results